how many liters of a 25% alcohol solution must be mixed wuth a 12% solution to get 13 liters of 15% solution.
Answer by Maths68(1474)

Solution A
Concentration = 25%=0.25
Amount = x
=======================
Solution B
Concentration = 12%=0.12
Amount = 13-x
==========================
Resultant Solution
Concentration = 15%=0.15
Amount = 13 liters
========================================
(Concentration of A * Amount of A)+(Concentration of B * Amount of B)=(Concentration of Resultant * Amount of Resultant)
==========================================
0.25x+0.12(13-x)=0.15*13
0.25x+1.56-0.12x=1.95
0.25x+1.56-0.12x=1.95
0.13x=1.95-1.56
0.13x=0.39
0.13x/13=0.39/13
x=3
==========================================
Amount of solution A = x = 3 liters
Amount of solution B = 13-x = 13-3 = 10 liters
Answer
=======
3 liters of a 25% alcohol solution must be mixed with 10 liters of 12% solution to get 13 liters of 15% solution.